A vehicle for landing on the moon has a mass of 4500kg. The damped spring-undercarriage system of the vehicle has a stiffness of 450kn/m and a damping ratio of 0.2. The rocket lift engines are cut off when the vehicle is hovering at an altitude of 10m. Calculate the maximum deflection of the undercarriage system when the vehicle hits the surface. On the moon, the gravitational acceleration is 1.6m/s2
1
Expert's answer
2012-09-25T11:45:17-0400
G=1.6m/s2
M=4500
F=MG
H=10m
A=MGh=4500*1.6*10= 450kn/m *xm*0.2
X=160/0.2=800